OpenStudy (anonymous):

Write the Hindu-Arabic numeral in expanded form. 703

Parth (parthkohli):

\[\overline{abc} = 100a + 10b + c\]

Parth (parthkohli):

So, for example, 612 = 600 + 10 + 2

OpenStudy (anonymous):

\[\left( 7\times10^{2} \right) + \left( 0\times10^{1} \right) + \left( 3\times 10^{0} \right)\]

OpenStudy (anonymous):

thats what I got

OpenStudy (anonymous):

i think middle term is wrong

OpenStudy (anonymous):

oh no it is right

OpenStudy (anonymous):

I was about to say please no :'(

OpenStudy (anonymous):

lol sorry it is right

OpenStudy (anonymous):

your right